OpenStudy (anonymous):

Find the number that must divide each term in the equation so that the equation can be solved by the method of completing the square: 6x^2-9x=19 a. 6 b. -9 c. x d. 19

OpenStudy (anonymous):

Thanks @wolf1728, I think the answer is 6.

OpenStudy (anonymous):

Is that right? I think it's 6 but I'm not sure!

OpenStudy (wolf1728):

6x^2-9x=19 To complete the square: Keep all terms to the left and move the "non X" term to the right: 6x^2 -9x = 19 Divide the equation by the coefficient of X² which in this case is 6 x² -1.5x = 3.16666666 Now for "completing the square" we: take the coefficient of X which is -1.5 divide that by 2 -.75 square that number .5625 then add it to both sides of the equation x² -1.5x +.5625= 3.16666666 +.5625 Take the square root of BOTH sides of the equation x-.75 = 1.931105036 x = 2.681105036 x-.75 = -1.931105036 x = -1.181105036 Well I completed the square but I don't know which of the options to choose. x +.75= 1.931105036

OpenStudy (anonymous):

Thank you so much!!!!!!

OpenStudy (wolf1728):

Sheesh it's just that simple? Divide by 6? Well, glad to help out. :-)

OpenStudy (anonymous):

To complete the square \[ (x+p)^2 = x^2 + 2px + p^2\] You have \[6x^2 - 9x =19\] ****You have to get rid of the coefficient in front of x to fit the form, so divide by 6**** and add p^2 to each side \[x^2 + \left( -\frac{ 9}{6}x \right) + (p^2) =-\frac{19}{6} +( p^2 )\] Solve for p \[ 2p = -\frac{9}{6} \] \[ p = -\frac{9}{12} = -\frac{3}{4} \] \[ p^2 = \frac{9}{16} \] \[(x-\frac{3}{4})^2 = x^2 -\frac{9}{12}x + \left( \frac{9}{16} \right) = \frac{19}{6} + \left( \frac{9}{16} \right) \] \[(x-\frac{3}{4})^2 = \frac{179}{48} \] \[x = \frac{3}{4} ± \sqrt{ \frac {179}{48}}\]